# Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels

### v1

@@ B.3 Integration with Vision-Language Models
In Table 8, we present an extension of the results in Table 3 in the main paper, including the standard deviation. Interestingly, the standard deviation for the reward obtained by using LICORICE with GPT-4o as the annotator does not always follow the same trend as shown in Table 7 (when we assume access to a more accurate human annotator). Instead, the standard deviation is relatively consistent for PixelCartPole-v1, regardless of the budget. It steadily decreases for DoorKey-7x7, as expected. However, in DynamicObstacles-5x5, we see an increase when $B=300$. We are not sure of the cause of this. Perhaps at this point the algorithm begins overfitting to the errors in the labels from GPT-4o (the concept error rate is the same for $200$ and $300$ labels). Further investigation is required to understand the underlying factors contributing to this anomaly.
Figure 4: Concept policy model trained with concept labels provided by GPT-4o waits until the coast is clear to move to the goal. It appears to make a small mistake in the bottom left, requiring it to wait slightly longer than necessary to navigate to the goal.
GPT-4o Concept Labeling Errors Table 5 list detailed concept errors for concepts in all environments. In PixelCartPole-v1, cart position and pole angle have smaller errors, while velocities require understanding multiple frames and thus are harder to predict accurately. For DoorKey-7x7 and DynamicObstacles-5x5, different concepts have slightly varying concept errors, indicating visual tasks have different difficulties for GPT-4o. Agent direction has as high as around 0.3 prediction error for both DoorKey-7x7 and DynamicObstacles-5x5. Direction movable is also hard for DoorKey-7x7, with a high concept error even if it is a binary concept. We posit it requires the correct understanding of more than one particular object to ensure correctness. The concept accuracies in DoorKey-7x7 are generally higher than DynamicObstacles-5x5, suggesting GPT-4o more struggles with a larger grid.
Example Behavior of Agent with GPT-4o-labeled concepts In Figure 4, we show an example of a GPT-4o-trained RL agent on DynamicObstacles-5x5, in which the agent appears to wait until it is safe to move towards the goal: the green square. The image sequence shows the agent (red triangle) starting from its initial position and moving to the right. It then is cornered by an obstacle (blue circle), then both obstacles. In the bottom left corner frame, it appears to make a mistake by turning to the right, meaning it missed a window to escape. Finally, it moves to the goal when the path is clear in the second-to-last and last frames (bottom right). This behavior highlights that the agent may still learn reasonable behavior even if the concept labels may be incorrect (causing it to make a mistake, as in the bottom left frame).
@@ B.4 Ablation
PixelCartPole-v1 LICORICE-IT 0.533 896±0.257 169plus-or-minus0.5338960.257169\$0.533\,896$\pm$\$0.257\,169\$0.533 896 ± 0.257 169 0.079 622±0.026 734plus-or-minus0.0796220.026734\$0.079\,622$\pm$\$0.026\,734\$0.079 622 ± 0.026 734
@@
DoorKey-7x7 LICORICE-IT 0.992 412±0.007 62plus-or-minus0.9924120.00762\$0.992\,412$\pm$\$0.007\,62\$0.992 412 ± 0.007 62 0.089 535±0.015 02plus-or-minus0.0895350.01502\$0.089\,535$\pm$\$0.015\,02\$0.089 535 ± 0.015 02
@@
DynamicObstacles-5x5 LICORICE-IT 0.996 889±0.002 139plus-or-minus0.9968890.002139\$0.996\,889$\pm$\$0.002\,139\$0.996 889 ± 0.002 139 0.001 202±0.000 982plus-or-minus0.0012020.000982\$0.001\,202$\pm$\$0.000\,982\$0.001 202 ± 0.000 982
@@
Table 9: Ablation study results for LICORICE in all environments, where we delete every single component to compare with LICORICE. This shows a more complete version of the results in Table 4.
@@
Figure 5: Learning curves for all ablations. Shaded region shows 95% CI, calculated using 1000 bootstrap samples.
@@
In Table 9, we present an extension of the results in Table 4, including standard deviation. Figure 5 shows all learning curves for ablations in all environments. In PixelCartPole-v1, we clearly see the benefit of iterative strategies on reward: LICORICE, LICORICE-DE and LICORICE-AC consistently increase in reward and converge at high levels, but LICORICE-IT converges at less than $50\%$ of the optimal reward. We also see that LICORICE-IT also achieves higher concept error, indicating that it struggles to learn the larger distribution of concepts induced by a non-optimal policy. In DoorKey-7x7, all algorithms steadily increase in reward. However, we can see a dip at around $10^{6}$ environment steps where we begin the second iteration for LICORICE, LICORICE-AC, and LICORICE-DE. Although LICORICE-IT achieves similar reward to LICORICE, LICORICE achieves lower concept error, which is beneficial for the goal of interpretability. Finally, in DynamicObstacles-5x5, LICORICE-AC lags behind the most in terms of both reward and concept error. This result indicates that the active learning component is most important for this environment.
Figure 6: Performance of all ablations at the final training iteration, with black bars representing the 95% confidence interval, calculated using 1000 bootstrap samples.
We now inspect the performance at the last training iteration in Figure 6. Overall, we find that LICORICE performs better than or equal to the ablations on all environments. It enjoys the best reward performance on PixelCartPole-v1, while performing similarly to LICORICE-IT on DoorKey-7x7 and to LICORICE-IT and LICORICE-DE on DynamicObstacles-5x5. It exhibits the lowest concept error on PixelCartPole-v1 and DoorKey-7x7, and achieves similarly low error to LICORICE-DE and LICORICE-IT on DynamicObstacles-5x5.
In PixelCartPole-v1, LICORICE achieves the highest reward, indicating the benefits of multiple iterations, active learning, and decorrelation, which are absent in LICORICE-IT, LICORICE-AC, and LICORICE-DE, respectively. LICORICE also enjoys the lowest concept error, showcasing its ability to learn and apply concepts accurately. LICORICE-AC and LICORICE-DE perform well in minimizing concept error, while LICORICE-IT shows a relatively higher error rate, underscoring the importance of multiple iterations in some environments.
@@ C.1 Limitations and Future Work
While our approach has demonstrated promising results, there are several limitations to be addressed in future work. One significant challenge is the difficulty VLMs face with certain types of concepts, especially continuous variables. This limitation can impact the overall performance of concept-based models, especially in domains where continuous data is prevalent. Addressing this issue could involve developing specialized techniques or using existing tools and libraries to better complement VLM capabilities.
Another area for future improvement is the refinement of our active learning and sampling schemes. Our current method employs an disagreement-based acquisition function to select the most informative data points for labeling. While this approach is effective, there is potential for exploring more sophisticated active learning strategies, such as incorporating advanced exploration-exploitation trade-offs or leveraging recent advancements in active learning algorithms [28].
Finally, designing a concept-based representation for RL remains an open challenge. Our work provides a few illustrative examples, but the exact design of these representations can significantly impact performance, often for reasons that are not entirely clear — especially when using VLMs as annotators. Prior work [7] proposed some desiderata for concepts in RL, but future work could refine these principles, especially in the face of VLM annotators. Future work could also include systematically investigating the factors that influence the effectiveness of different concept-based representations in RL. This could involve extensive empirical studies, theoretical analyses, and the development of new design principles that guide the creation of effective concept representations. Understanding these factors better will help in creating more reliable and interpretable RL models, ultimately advancing the field and broadening the applicability of concept-based approaches in various RL tasks.
@@ C.2 Broader Impacts
Interpretability in RL
Incorporating concept learning with RL presents both positive and negative societal impacts. On the positive side, promoting interpretability and transparency in decision-making fosters trust and accountability. However, in cases where it yields incorrect results, stakeholders might be misled into trusting flawed decisions due to the perceived transparency of the model [12]. Unintended misuse could also occur if stakeholders lack the technical expertise to accurately interpret the models, leading to erroneous conclusions and potentially harmful outcomes. To mitigate these risks, an avenue for future work is developing clear guidelines for interpreting these models and tools to scaffold non-experts’ understanding of the model outputs.
Using VLMs for Concept Labeling

### v2

@@ B.3 Integration with Vision-Language Models
In Table 7, we present an extension of the results in Figure 3 in the main paper, including the standard deviation. Interestingly, the standard deviation for the reward obtained by using LICORICE with GPT-4o as the annotator does not always follow the same trend as shown in Table 6 (when we assume access to a more accurate human annotator). Instead, the standard deviation is relatively consistent for PixelCartPole, regardless of the budget. It steadily decreases for DoorKey, as expected. However, in DynamicObstacles, we see an increase when $B=300$. We are not sure of the cause of this. Perhaps at this point the algorithm begins overfitting to the errors in the labels from GPT-4o (the concept error rate is the same for $200$ and $300$ labels). Further investigation is required to understand the underlying factors contributing to this anomaly.
Figure 8: LICORICE+GPT-4o waits until the coast is clear to move to the goal. It appears to make a small mistake in the bottom left, requiring it to wait slightly longer than necessary to navigate to the goal.
Different tasks and concepts have various difficulties for GPT-4o. Table 4 list detailed concept errors for concepts in all environments. In PixelCartPole, cart position and pole angle have smaller errors, while velocities require understanding multiple frames and thus are harder to predict accurately. For DoorKey and DynamicObstacles, different concepts have slightly varying concept errors, indicating visual tasks have different difficulties for GPT-4o. Agent direction has as high as around 0.3 prediction error for both DoorKey and DynamicObstacles. Direction movable is also hard for DoorKey, with a high concept error even if it is a binary concept. We posit it requires the correct understanding of more than one particular object to ensure correctness. The concept accuracies in DoorKey are generally higher than DynamicObstacles, suggesting GPT-4o more struggles with a larger grid.
A GPT-4o-trained RL agent recovers from a mistake. In Figure 8, we show an example of a GPT-4o-trained RL agent on DynamicObstacles, in which the agent appears to wait until it is safe to move towards the goal: the green square. The image sequence shows the agent (red triangle) starting from its initial position and moving to the right. It then is cornered by an obstacle (blue circle), then both obstacles. In the bottom left corner frame, it appears to make a mistake by turning to the right, meaning it missed a window to escape. Finally, it moves to the goal when the path is clear in the second-to-last and last frames (bottom right). This behavior highlights that the agent may still learn reasonable behavior even if the concept labels may be incorrect (causing it to make a mistake, as in the bottom left frame).
@@ B.4 Ablation
PixelCartPole LICORICE-IT 0.533 896±0.257 169plus-or-minus0.5338960.257169\$0.533\,896$\pm$\$0.257\,169\$0.533 896 ± 0.257 169 0.079 622±0.026 734plus-or-minus0.0796220.026734\$0.079\,622$\pm$\$0.026\,734\$0.079 622 ± 0.026 734
@@
DoorKey LICORICE-IT 0.992 412±0.007 62plus-or-minus0.9924120.00762\$0.992\,412$\pm$\$0.007\,62\$0.992 412 ± 0.007 62 0.089 535±0.015 02plus-or-minus0.0895350.01502\$0.089\,535$\pm$\$0.015\,02\$0.089 535 ± 0.015 02
@@
DynamicObstacles LICORICE-IT 0.996 889±0.002 139plus-or-minus0.9968890.002139\$0.996\,889$\pm$\$0.002\,139\$0.996 889 ± 0.002 139 0.001 202±0.000 982plus-or-minus0.0012020.000982\$0.001\,202$\pm$\$0.000\,982\$0.001 202 ± 0.000 982
@@
Boxing LICORICE-IT 0.978 152±0.036 465plus-or-minus0.9781520.036465\$0.978\,152$\pm$\$0.036\,465\$0.978 152 ± 0.036 465 0.135 606±0.069 144plus-or-minus0.1356060.069144\$0.135\,606$\pm$\$0.069\,144\$0.135 606 ± 0.069 144
LICORICE-DE 1.001 046±0.042 38plus-or-minus1.0010460.04238\$1.001\,046$\pm$\$0.042\,38\$1.001 046 ± 0.042 38 0.044 79±0.023 005plus-or-minus0.044790.023005\$0.044\,79$\pm$\$0.023\,005\$0.044 79 ± 0.023 005
LICORICE-AC 0.994 306±0.009 63plus-or-minus0.9943060.00963\$0.994\,306$\pm$\$0.009\,63\$0.994 306 ± 0.009 63 0.087 981±0.063 943plus-or-minus0.0879810.063943\$0.087\,981$\pm$\$0.063\,943\$0.087 981 ± 0.063 943
LICORICE 1.004 997±0.017 713plus-or-minus1.0049970.017713\$1.004\,997$\pm$\$0.017\,713\$1.004 997 ± 0.017 713 0.049±0.010 967plus-or-minus0.0490.010967\$0.049$\pm$\$0.010\,967\$0.049 ± 0.010 967
@@
Table 9: Ablation study results for LICORICE in the first four environments, where we study the effect of each component to compare with LICORICE. This shows a more complete version of the results in Table 3.
@@
Figure 9: Learning curves for ablations in the first three environments. Shaded region shows 95% CI, calculated using 1000 bootstrap samples.
@@
In Table 9, we present an extension of the results in Table 3, including standard deviation. Figure 9 shows all learning curves for ablations in all environments. In PixelCartPole, we clearly see the benefit of iterative strategies on reward: LICORICE, LICORICE-DE and LICORICE-AC consistently increase in reward and converge at high levels, but LICORICE-IT converges at less than $50\%$ of the optimal reward. We also see that LICORICE-IT also achieves higher concept error, indicating that it struggles to learn the larger distribution of concepts induced by a non-optimal policy. In DoorKey, all algorithms steadily increase in reward. However, we can see a dip at around $10^{6}$ environment steps where we begin the second iteration for LICORICE, LICORICE-AC, and LICORICE-DE. Although LICORICE-IT achieves similar reward to LICORICE, LICORICE achieves lower concept error, which is beneficial for the goal of interpretability. Finally, in DynamicObstacles, LICORICE-AC lags behind the most in terms of both reward and concept error. This result indicates that the active learning component is most important for this environment.
Impact on data diversity. We hypothesize that data decorrelation enables better performance by increasing the diversity of the training data. As a result, we investigate the role of data decorrelation on state diversity by calculating the average distance to the $k$ nearest samples in the concept label space for all labeled samples. The intuition is that a larger average distance to nearest neighbors indicates samples are more spread out in the space, suggesting higher diversity in the training data. Let $D=\{(s_{1},c_{1}),\dots,(s_{b},c_{b})\}$ be the labeled dataset and $\text{Neighbor}(c_{i},k)$ be the index set of $k$ nearest neighbors with $k=10$ for all environments, then we calculate diversity as:
$d=\frac{1}{bk}\sum_{i=1}^{b}\sum_{j\in\text{Neighbor}(c_{i},k)}||c_{i}-c_{j}||.$
We present $d$ along with the reward in Table 10 for LICORICE and LICORICE-DE. As hypothesized, decorrelation increases diversity in 3 of the 4 environments. LICORICE achieves higher reward than LICORICE-DE. In DynamicObstacles, the state diversity and reward is similar for both LICORICE and LICORICE-DE. We believe that, in this environment, decorrelation is less critical because neighboring states are inherently more dissimilar due to the dynamic nature of the obstacles.
LICORICE LICORICE-DE
Environment $d$ $R$ $d$ $R$
PixelCartPole 0.12 $\pm$ 0.00 $1.0$\pm$0.0$$ 0.10 $\pm$ 0.01 0.9669±0.051 17plus-or-minus0.96690.05117\$0.9669$\pm$\$0.051\,17\$0.9669 ± 0.051 17
DoorKey 1.12 $\pm$ 0.18 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.25 $\pm$ 0.00 0.783 113±0.079 936plus-or-minus0.7831130.079936\$0.783\,113$\pm$\$0.079\,936\$0.783 113 ± 0.079 936
DynamicObstacles 0.93 $\pm$ 0.04 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.99 $\pm$ 0.09 0.980 537±0.008 092plus-or-minus0.9805370.008092\$0.980\,537$\pm$\$0.008\,092\$0.980 537 ± 0.008 092
Boxing 7.20 $\pm$ 0.20 0.968 879±0.036 656plus-or-minus0.9688790.036656\$0.968\,879$\pm$\$0.036\,656\$0.968 879 ± 0.036 656 6.53 $\pm$ 0.08 0.970 273±0.009 912plus-or-minus0.9702730.009912\$0.970\,273$\pm$\$0.009\,912\$0.970 273 ± 0.009 912
Table 10: Training data diversity ($d$) and reward ($R$) with and without data decorrelation. Higher $d$ is better (more diverse). For each environment, the highest $d$ is bolded. When the standard deviations overlap, we bold both.
B.5 Hyperparameter Sensitivity
Figure 10: Hyperparameter sensitivity results for DynamicObstacles. Bars represent standard deviation. Here, bars are grouped by the same $p$ values (acceptance probability). Bars of the same color share the same $\tau$ value (budget multiplier), and bars with the same pattern share the same value for $N$ (committee size). Generally, $p=0.1$ tend to perform better overall with greater stability. Having a larger committee ($N=5$ vs. $N=3$) also tends to help with performance.
We now seek to understand how sensitive LICORICE is to the choice of hyperparameters. Figure 10 shows the results of this experiment for DynamicObstacles. Interestingly, we find that the main difference between the settings is the variation. Overall, LICORICE is relatively robust to the choice of hyperparameter settings in this environment, with the final average reward ranging between 96% and 99.5% of the optimal. The concept error for all settings was $00.00\pm 0$ 0.00, so we do not plot it here.
B.6 Choice of Acquisition Strategy
Budget 100 200 300
R c Error R c Error R c Error
Disagreement 0.96 0.05 0.97 0.01 1.00 0.00
Entropy 0.95 0.07 0.98 0.02 0.99 0.00
Table 11: Comparison of disagreement- and entropy-based active learning strategies for DynamicObstacles under different labeling budgets. There is essentially no difference between the two strategies.
We now investigate the impact of the choice of acquisition strategy on both reward and concept accuracy. We implemented an entropy-based disagreement measurement for the committee and studying its effects on a MiniGrid environment. Specifically, we implement the following vote entropy acquisition function: $U(s_{i})=-\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{P_{k}}\frac{V_{ijk}}{N}\log\frac{V_{ijk}}{N}$, where $s_{i}$ is the input state, $K$ is the number of concepts we learn, $P_{k}$ is the number of classes for the $k$-th concept, $V_{ijk}$ is the number of votes for the $j$-th class of the $k$-th concept for state $s_{i}$, and $N$ is the number of committee members. Table 11 shows these results on DynamicObstacles. There is essentially no difference between the two strategies in terms of both reward and concept error.
@@ C.1 Limitations and Future Work
While our approach has demonstrated promising results, there are several limitations to be addressed in future work. One significant challenge is the difficulty VLMs face with certain types of concepts, especially continuous variables. This limitation can impact the overall performance of concept-based models, especially in domains where continuous data is prevalent. Furthermore, although VLMs can be successfully used for automatic labeling of some concepts, there are still hallucination issues (Achiam et al., 2023) and other failure cases, such as providing inaccurate counts. We believe that future work that seeks to improve general VLM capabilities and mitigate hallucinations would also help overcome this limitation. Addressing this issue could involve developing specialized techniques or using existing tools and libraries to better complement VLM capabilities.
Another area for future improvement is the refinement of our active learning and sampling schemes. Our current method employs an disagreement-based acquisition function to select the most informative data points for labeling. While this approach is effective, there is potential for exploring more sophisticated active learning strategies, such as incorporating advanced exploration-exploitation trade-offs or leveraging recent advancements in active learning algorithms (Tharwat & Schenck, 2023).
Furthermore, scaling challenges may arise from environmental complexity, such as when only a subset of given concepts are relevant or when learning certain concepts is prerequisite to gathering training data for others. Future research could explore using attention mechanisms (Vaswani et al., 2017) or sparse coding (Olshausen & Field, 1996) to identify relevant concepts. Another exciting direction is the work in human-AI complementarity and learning-to-defer algorithms (Mozannar et al., 2023; Verma et al., 2023) to train an additional classifier for deferring labeling to a person when the chance of an error is high. In our work, we assume a known set of concepts for the environment; future work could investigate the use of human-VLM teams to determine a reasonable concept set for the environment.
Finally, designing a concept-based representation for RL remains an open challenge. Our work provides a few illustrative examples, but the exact design of these representations can significantly impact performance, often for reasons that are not entirely clear — especially when using VLMs as annotators. Prior work (Das et al., 2023) proposed some desiderata for concepts in RL, but future work could refine these principles, especially in the face of VLM annotators. Future work could also include systematically investigating the factors that influence the effectiveness of different concept-based representations in RL. This could involve extensive empirical studies, theoretical analyses, and the development of new design principles that guide the creation of effective concept representations. Understanding these factors better will help in creating more reliable and interpretable RL models, ultimately advancing the field and broadening the applicability of concept-based approaches in various RL tasks.
@@ C.2 Broader Impacts
Interpretability in RL.
Incorporating concept learning with RL presents both positive and negative societal impacts. On the positive side, promoting interpretability and transparency in decision-making fosters trust and accountability. However, in cases where it yields incorrect results, stakeholders might be misled into trusting flawed decisions due to the perceived transparency of the model (Kaur et al., 2020). Unintended misuse could also occur if stakeholders lack the technical expertise to accurately interpret the models, leading to erroneous conclusions and potentially harmful outcomes. To mitigate these risks, an avenue for future work is developing clear guidelines for interpreting these models and tools to scaffold non-experts’ understanding of the model outputs.
Using VLMs for concept labeling.